OpenStudy (anonymous):

A line passes through the point (0, –4) and has a slope of . What is the equation of the line?

OpenStudy (e.mccormick):

Know your point-slope formula?

OpenStudy (anonymous):

I really just need the answer so i don't fail summer school...

OpenStudy (anonymous):

y-y1=m(x-x1) just plugin your values

OpenStudy (e.mccormick):

Then put it in \(y-y_1=m(x-x_1)\) and get an answer.

OpenStudy (anonymous):

what is the slope of the line?

OpenStudy (anonymous):

so if it has a slope of 3/2 where do i plug that in

OpenStudy (anonymous):

x1=0 y1=-4 y-(-4)=m(x-0) y+4=mx y=-4+(3/2)x

OpenStudy (anonymous):

what

OpenStudy (anonymous):

the equation of the line is y=-4+(3/2)x plugin in (3/2) in place of m

OpenStudy (anonymous):

can you just tell me the answer dude please

OpenStudy (anonymous):

Which equation is equivalent to 3x – 4y = 8? A. y=3/4x+2 B.y=-3/4x+2 C.y=-3/4x-2 D.y=3/4x-2 and this one its hard

OpenStudy (anonymous):

im in special ed...

OpenStudy (anonymous):

the answer is D